I'm asked to determine a quadratic equation with the information given and solve.
The perimeter of a rectangle is 34cm and the diagonal is 13cm.
Determine the lengths of the sides.



Answer by Alan3354(69443)

The perimeter of a rectangle is 34cm and the diagonal is 13cm.
--------
P = 2W + 2L = 34
W + L = 17 --> L = 17-W
----
W^2 + (17-W)^2 = 13^2
Solve for W, then for L.